OpenStudy (anonymous):

What is the solution to the equation? (-9x+5)^1/2=(3+4x)^1/2

OpenStudy (anonymous):

\[\sqrt{(-9x+5)}=\sqrt{(3+4x)}\]

OpenStudy (anonymous):

Square both the sides first.. \[-9x + 5 = 3 + 4x \]

OpenStudy (anonymous):

Add \(9x\) to both the sides: \(5 = 3 + 13 x\) now, subtract \(3\) from both the sides: \(2 = 13x\)

OpenStudy (anonymous):

Now divide by \(13\) both the sides to find \(x\)..

OpenStudy (anonymous):

2/13=x

OpenStudy (anonymous):

yep.. Good.. :)

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

\(\dagger\)